# Software-defined lock-in demodulator for low-frequency resistance noise measurements

### v1

@@ Abstract
The following article has been submitted to Review of Scientific Instruments. After it is published, it will be found at https://pubs.aip.org/aip/rsi.
The resolution of low-frequency resistance noise measurements is increased by amplitude modulation, shifting the spectrum of the resistance fluctuations away from the noise contributed by measurement instruments. However, commercial lock-in amplifiers used for de-modulating the fluctuations exhibit a problematic noise contribution, which imposes a hard lower limit on the relative resistance noise that can be detected. We replace the lock-in amplifier hardware by equivalent digital signal processing performed using open-source software and inexpensive data acquisition systems. Our solution offers superior low-frequency noise performance with a reduction of the voltage power spectral density by about two orders of magnitude compared to high-end commercial instruments.
@@ I Introduction
of this background has a value of for the Stanford Research Systems (SRS) SR830 instrument commonly found in physics laboratories. The quadratic dependence of this background on the mean voltage matches that of the DUT noise voltage drop, for which according to Ohm’s law, making them difficult to distinguish. Therefore, the measurement resolution is limited if a bridge configuration Scandurra et al. (2022) to suppress is not available. To circumvent this problem, we propose to replace the lock-in amplifier hardware with software digital signal processing. In the following, the design of this software solution, as well as the accompanying hardware is described, and a significant improvement in noise performance is demonstrated.
@@ II Implementation
Fig. 1a shows the circuit used for noise measurements with conventional lock-in amplifier hardware, in which a function generator (e.g. SRS DS360) with low harmonic distortion excites the DUT resistance with a current set by . The DUT voltage drop is amplified and passed into the lock-in instrument, which includes a mixer and a low-pass filter, and which outputs the demodulated noise signal to be recorded by an analogue-to-digital converter (ADC). The noise power spectral density is then calculated from the recorded time series. For the software-based technique, see Fig. 1b, the DUT voltage is buffered by a low-noise, multi-stage amplifier (SRS SR560), whose output is directly recorded by an ADC. A second, simultaneously sampling ADC records the function generator output , which will serve as a reference for the demodulation conducted in software. The differential, isolated ADCs of the Delta-Sigma-type Kester (2005) are provided by a data acquisition system (DAQ, NI 9239), which features a high resolution (), sufficient bandwidth () and a low noise floor () with an input range of .
@@
The software signal processing system, which we make available under an open-source license at Thyzel (2024), is visualized in Fig. 2. Its acquisition stage employs driver modules, in this case the NI DAQmx framework via the nidaqmx Python package, to set up and read data from the ADCs. Instead of processing samples of the input voltage one-by-one, double buffering allows an uninterrupted, block-wise readout. This is advantageous because, although the execution speed of Python is limited by the interpreter, a large ecosystem of high-quality scientific libraries (NumPy Harris et al. (2020), SciPy Virtanen et al. (2020)) is available that deal efficiently with arrayed data, such as time series blocks. Block-wise signal processing is implemented using the Observer software architecture pattern Gamma, Helm, and Vlissides (1995), also known as Publish-Subscribe. Once the acquisition driver, an Observable object, has filled one of its two buffers, it emits the buffer contents as a time series block, which is received by one or more signal processing Observers. Each Observer object applies a transformation (e.g. filtering, mixing, downsampling, …) to the time series data in the block and emits the result to other Observers further along the chain. In this manner, a quasi-real-time signal processing pipeline is built by composing Observable and Observer objects.
@@
At the end of the acquisition stage, the first signal processing step applies a low-pass filter on both channels independently to avoid aliasing in the mixing steps to follow. Butterworth-type filters of sixth order are used, which offer an ideally flat passband important for broad-band noise measurements, and a good roll-off rate of per octave Oppenheim and Schafer (1975). Their infinite impulse response allows an implementation by a difference equation, which offers higher computational efficiency than filters based on convolution, and which is available through a combination of the functions scipy.signal.butter and scipy.signal.sosfilt. Filter state is retained across time series blocks using the zi argument.
@@
The time series blocks produced by the demodulator stage are passed to the output stage, where they are down-sampled, or decimated, to reduce the data rate. The new sampling rate is twice the demodulator cut-off frequency, plus some oversampling to account for finite filter attenuation in the stopband. After downsampling, the time series blocks are concatenated and written to a compressed binary file in the HDF5 format The HDF Group (2024) for later analysis. Thanks to the streaming software architecture, the output signal can also be monitored graphically in quasi-real-time. For this purpose, the time series blocks are buffered and re-sized for plotting in a graphical interface based on the Matplotlib library Hunter (2007). Also, the noise power spectral density is computed by a standard Fourier periodogram using the scipy.signal.welch method and displayed in real time, giving immediate feedback on the noise properties of the DUT.
@@ III Performance tests
Having verified the frequency response, we next measured the background noise of both the SR830 hardware and the software lock-in implementation. For this purpose, a noise-free device under test (DUT) was simulated by connecting the function generator DS360 directly to the lock-in inputs. Fig. 3b shows that SR830 exhibits intrinsic -type noise with , whose magnitude depends quadratically on the input amplitude . This makes it difficult to distinguish from noise originating from a DUT, which has the same scaling behaviour. It also makes it impossible, without resorting to cross-correlation or bridge circuits suppressing Thyzel et al. (2024), to measure any DUT resistance noise smaller than . In contrast, the input noise of the software-defined lock-in setup shown in Fig. 3c does not exhibit any such dependence on the input amplitude. At an input amplitude of , a significant reduction in background noise power density of about two orders of magnitude at is achieved. This means that by using multi-stage amplifiers to exploit the higher input range of for NI 9239, the software-based lock-in method can offer superior low-frequency noise performance compared to the SR830 hardware instrument.
Finally, we investigated the noise background of the software lock-in implementation at very low frequencies in more detail. For this purpose, we again fed a sine wave directly into the NI 9239 channels. Three input channels were demodulated independently, but sharing the same reference channel. The noise spectrum shown in Figure 4 exhibits an -like upturn at extremely low frequencies corresponding to a time scale of about three seconds and longer. Surprisingly, this part of the spectrum seems to be fully correlated between the three input channels, as the cross power spectral density Sampietro, Fasoli, and Ferrari (1999) computed by scipy.signal.csd matches the single-channel spectrum. This indicates that the contribution is caused either by the common reference channel, or by the sine signal source. We speculate that in the latter case, noise in the oscillator frequency driving the signal generator output could cause a -like modulation of the sine signal. In any case, the noise magnitude is only approximately , two orders of magnitude lower than the background of the SR830 instrument.
@@ IV Conclusion and outlook
We have implemented a software-defined lock-in amplifier for use in amplitude-modulated low-frequency resistance noise measurements. Our streaming, quasi-real-time software architecture allows for highly efficient signal processing utilizing the well-established scientific Python library ecosystem and relatively inexpensive data acquisition hardware. We showed the correctness of the software implementation by comparing its output noise spectrum with that of a commercial, high-end lock-in amplifier instrument. Importantly, our software implementation exhibits a noise power spectral density about two orders of magnitude below that of the commercial instrument, increasing the resolution for low-noise samples significantly when bridge setups are not an option. Future improvements could include scaling up the number of DAQ channels and pre-amplifiers in a cross-correlation setup to suppress the remaining noise background further, as suggested in Ref. Scandurra, Giusi, and Ciofi (2013). Furthermore, multiple reference channels, instead of a single one, could possibly be used to reduce the residual contribution. Aside from applications in noise spectroscopy, a true phase-sensitive detection using a software-based phase-locked loop would allow our low-cost solution to be used for complex impedance spectroscopy at low frequencies.

### v2

@@ Abstract
The resolution of low-frequency resistance noise measurements can be increased by amplitude modulation, shifting the spectrum of the resistance fluctuations away from the noise contributed by measurement instruments. However, commercial lock-in amplifiers used for de-modulating the fluctuations exhibit a problematic noise contribution, which imposes a hard lower limit on the relative resistance noise that can be detected. We replace the lock-in amplifier hardware by equivalent digital signal processing performed using open-source software and off-the-shelf data acquisition systems. Compared to previous implementations of the lock-in principle, our solution offers real-time preview capabilities and is resource-efficient for long acquisition times at high sampling rates. Importantly, compared to high-end commercial lock-in instruments, our system offers superior low-frequency noise performance with a reduction of the voltage power spectral density by about two orders of magnitude.
@@ I Introduction
of this background has a value of for the Stanford Research Systems (SRS) SR830 instrument commonly found in physics laboratories. The quadratic dependence of this background on the mean voltage matches that of the DUT noise voltage drop, for which according to Ohm’s law, making them difficult to distinguish. Therefore, the measurement resolution is limited if a bridge configuration Scandurra et al. (2022) to suppress is not available.
To circumvent this problem, we replace the lock-in amplifier hardware with software digital signal processing. Our implementation, specifically designed for low-frequency noise measurements, offers advantages in ease of use and resource efficiency over previous reportsUhl, Bruder, and Stienkemeier (2021); O’Donoghue and Cantillon-Murphy (2015); Andersson et al. (2007). First and foremost, noise measurements usually necessitate the acquisition of signals over durations up to one hour. If demodulation takes place only after the full input signal has been acquired as in Ref. Uhl, Bruder, and Stienkemeier (2021), the output cannot be evaluated before the experiment ends. Therefore, we implement a streaming architecture where lock-in demodulation takes place as the input data is being acquired, allowing a quasi-real-time preview of the output signal. This feature has not been previously reported, see Fig. 1, and it makes our solution a drop-in replacement for conventional lock-in amplifier hardware.
Second, the lock-in carrier frequency, and therefore the alias-free sampling rate, is often far higher – to – than the frequency scales of interest for noise measurements – e.g. to . Storing the demodulated output signal at this sampling rate in memory and persistent storage as in Ref. Uhl, Bruder, and Stienkemeier (2021) is inefficient, producing data at a rate of per hour when sampling at . Therefore, the demodulator output should be downsampled/decimated as it is being acquired, leading to a lower output data rate and a smaller memory footprint. This is advantageous for integrating the lock-in amplifier into a larger signal-processing system, which may perform computationally expensive operations such as Fourier transforms on the demodulated output or archive it in long-term storage. A previous implementation achieves downsampling by conducting the signal processing on a microcontroller-based hardware system O’Donoghue and Cantillon-Murphy (2015), which is highly cost-effective, but requires the design and manufacturing of custom electronics. Our solution makes use of commercial, off-the-shelf data acquisition hardware and implements alias-free, streaming downsampling in software. Thanks to the careful use of buffers, the software does not require real-time scheduling, and therefore runs on general-purpose computer operating systems. Furthermore, a modular architecture allows the use of acquisition hardware from different vendors, an advantage over custom electronics in varying laboratory environments. In the following, the design of this open-source software package, as well as the accompanying hardware setup is described, and a significant improvement in low-frequency noise performance is demonstrated over lock-in amplifier hardware.
@@ II Implementation
Fig. 2a shows the circuit used for noise measurements with conventional lock-in amplifier hardware, in which a function generator (e.g. SRS DS360) with low harmonic distortion excites the DUT resistance with a current set by . The DUT voltage drop is amplified and passed into the lock-in instrument, which includes a mixer and a low-pass filter, and which outputs the demodulated noise signal to be recorded by an analogue-to-digital converter (ADC). The noise power spectral density is then calculated from the recorded time series. Note that this setup is only sufficient if the DUT resistance noise is not bias-dependent, which needs to be verified in all practical applications using suitable DC methods Scandurra et al. (2022); Thyzel et al. (2024). For the software-based technique, see Fig. 2b, the DUT voltage is buffered by a low-noise, multi-stage amplifier (SRS SR560), whose output is directly recorded by an ADC. A second, simultaneously sampling ADC records the function generator output , which will serve as a reference for the demodulation conducted in software. The differential, isolated ADCs of the Delta-Sigma-type Kester (2005) are provided by a data acquisition system (DAQ, NI 9239), which features a high resolution (), sufficient bandwidth () and a low noise floor () with an input range of .
@@
The software signal processing system, which we make available under an open-source license at Ref. Thyzel (2024), is visualized in Fig. 3. Its acquisition stage employs driver modules, in this case the NI DAQmx framework via the nidaqmx Python package, to set up and read data from the ADCs. Note that data acquisition hardware from other vendors can be easily integrated by implementing a new acquisition driver module. Instead of processing samples of the input voltage one-by-one, double buffering allows an uninterrupted, block-wise readout. This is advantageous because, although the execution speed of Python is limited by the interpreter, a large ecosystem of high-quality scientific libraries (NumPy Harris et al. (2020), SciPy Virtanen et al. (2020)) is available that deal efficiently with arrayed data, such as time series blocks. Furthermore, this buffered design is necessary on computer operating systems that do not support real-time scheduling, because it cannot be guaranteed that a single sample will be fully processed before the next one is acquired, as is the case in microcontroller-based implementations such as Ref. O’Donoghue and Cantillon-Murphy (2015). Block-wise signal processing is implemented using the Observer software architecture pattern Gamma, Helm, and Vlissides (1995), also known as Publish-Subscribe. Once the acquisition driver, an Observable object, has filled one of its two buffers, it emits the buffer contents as a time series block, which is received by one or more signal processing Observers. Each Observer object applies a transformation (e.g. filtering, mixing, downsampling, …) to the time series data in the block and emits the result to other Observers further along the chain. In this manner, a quasi-real-time signal processing pipeline is built by composing Observable and Observer objects.
@@
At the end of the acquisition stage, the first signal processing step applies a low-pass filter on both channels independently to avoid aliasing in the mixing steps to follow. Butterworth-type filters of sixth order are used, which offer an ideally flat passband important for broad-band noise measurements, and a good roll-off rate of per octave Oppenheim and Schafer (1975). Their infinite impulse response allows an implementation by a difference equation, which offers higher computational efficiency than filters based on convolution, and which is available through a combination of the functions scipy.signal.butter and scipy.signal.sosfilt. Filter state is retained across time series blocks using the zi argument, which is vital to avoid introducing discontinuities across block boundaries, as illustrated in Fig. 1.
@@
The time series blocks produced by the demodulator stage are passed to the output stage, where they are down-sampled, or decimated, to reduce the data rate. Care is taken to maintain correct sample spacing across block boundaries. The new sampling rate is twice the demodulator cut-off frequency, plus some oversampling to account for finite filter attenuation in the stopband. After downsampling, the time series blocks are concatenated and written to a compressed binary file in the HDF5 format The HDF Group (2024) for later analysis. Thanks to the streaming software architecture, the output signal can also be monitored graphically in quasi-real-time. For this purpose, the time series blocks are buffered and re-sized for plotting in a graphical interface based on the Matplotlib library Hunter (2007). Also, the noise power spectral density is computed by a standard Fourier periodogram using the scipy.signal.welch method and displayed in real time, giving immediate feedback on the noise properties of the DUT.
@@
Specific details about the software environment, installation and usage, as well as the implementation of new acquisition hardware are given in the README.md file of the software package, which is available for download at Ref. Thyzel (2024). The modular code can be integrated into existing signal processing algorithms, see the “Development” section of that file.
@@ III Performance tests
Having verified the frequency response, we next measured the background noise of both the SR830 hardware and the software lock-in implementation. For this purpose, a noise-free device under test (DUT) was simulated by connecting the function generator DS360 directly to the inputs of SR830 and NI 9239 in parallel. The lock-in amplifier instrument and the software-based solution therefore demodulated the same, un-amplified sinusoidal voltage signal with identical source and load impedances, and the same source noise level. The circuit remained unchanged in all measurements with both devices.
Fig. 4b shows that SR830 exhibits intrinsic -type noise with , whose magnitude depends quadratically on the input amplitude . This makes it difficult to distinguish from noise originating from a DUT, which has the same scaling behaviour. It also makes it impossible, without resorting to cross-correlation or bridge circuits suppressing Thyzel et al. (2024), to measure any DUT resistance noise smaller than . In contrast, the input noise of the software-defined lock-in setup shown in Fig. 4c does not exhibit any such dependence on the input amplitude. At an input amplitude of , a significant reduction in background noise power density of about two orders of magnitude at is achieved. In fact, the behavoiur of the software implementation resembles that of an ideal lock-in amplifier more closely, in that the input amplifier’s noise is frequency-shifted and therefore not visible. This is not the case for the SR830 instrument, whose output exhibits very strong noise. As the signal at the inputs of both instruments was identical, SR830’s undesirable noise, which we verified under various other measurement conditions, must be caused either by the digital algorithms in its demodulator or the semiconductor components in its amplifier stages and ADC. Note, however, that it is atypical for the input-referred noise level of simple amplifier or ADC circuits to increase with the input amplitude like in Fig. 4b. In conclusion, it remains unknown what causes the inconvenient noise in SR830, but our implementation of the lock-in method can offer superior low-frequency noise performance, in particular when multi-stage amplifiers are used to exploit the higher input range of for NI 9239.
Finally, we investigated the noise background of the software lock-in implementation at very low frequencies in more detail. For this purpose, we again fed a sine wave directly into the NI 9239 channels. Three input channels were demodulated independently, but sharing the same reference channel. The noise spectrum shown in Figure 5 exhibits an -like upturn at extremely low frequencies corresponding to a time scale of seconds. Surprisingly, this part of the spectrum seems to be fully correlated between the three input channels, as the cross power spectral density Sampietro, Fasoli, and Ferrari (1999) computed by scipy.signal.csd matches the single-channel spectrum. This indicates that the contribution is caused either by the common reference channel, or by the sine signal source. We speculate that in the latter case, noise in the oscillator frequency driving the signal generator output could cause a -like modulation of the sine signal. In any case, the noise magnitude is only approximately , two orders of magnitude lower than the background of the SR830 instrument.
@@ IV Conclusion and outlook
We have implemented a software-defined lock-in amplifier for use in amplitude-modulated low-frequency resistance noise measurements. Compared to previous implementations, our streaming software architecture allows for highly efficient signal processing with real-time preview capabilities, utilizing commercially available data acquisition hardware. We show the correctness of the software implementation by comparing its output noise spectrum with that of the commercial, high-end lock-in amplifier instrument Stanford Research Systems SR830. Importantly, our implementation exhibits a noise power spectral density at about two orders of magnitude below that of the commercial instrument. This makes it possible to investigate devices with an extremely low noise level, which could previously only be detected in bridge circuit configurations that are not available for all device types. Future improvements could include scaling up the number of DAQ channels and pre-amplifiers in a cross-correlation setup to suppress the remaining noise background further, as suggested in Ref. Scandurra, Giusi, and Ciofi (2013). Furthermore, multiple reference channels, instead of a single one, could possibly be used to reduce the residual contribution. Aside from applications in noise spectroscopy, a dual-phase detection using a software-based phase-locked loop as in Ref. Uhl, Bruder, and Stienkemeier (2021) would allow our solution to be used for complex impedance spectroscopy at low frequencies.